# Gemstones: A Model Suite for Multi-Faceted Scaling Laws

### v1

@@ Abstract
Scaling laws are typically fit using a family of models with a narrow range of frozen hyper-parameter choices. In this work we study scaling laws using a wide range of architecture and hyper-parameter choices, and highlight their impact on resulting prescriptions. As a primary artifact of our research, we release the Gemstones: the most comprehensive open-source scaling law dataset to date, consisting of over 4000 checkpoints from transformers with up to 2 billion parameters; these models have been trained with different learning rates, cooldown schedules, and architectural shapes. Our checkpoints enable more complex studies of scaling, such as a law that predicts language modeling performance as a function of model width and depth. By examining the various facets of our model suite, we find that the prescriptions of scaling laws can be highly sensitive to the experimental design process and the specific model checkpoints used during fitting.
@@ 1 Introduction
In this work, we produce a vast array of model checkpoints for studying how model design and model selection impact scaling laws. Our models, called the Gemstones because they are loosely based on scaled-down variants of the Gemma architecture, vary in their parameter count, width/depth ratio, training tokens, learning rates, and cooldown schedules. By fitting scaling laws to these checkpoints, we confirm that scaling law parameters and interpretations indeed depend strongly on the selection of models and fitting procedure used, and we quantify the degree to which these decisions impact predictions. By exploiting the variation among our model checkpoints, we also fit a number of unique scaling laws and analyze their predictions to discern whether they are consistent with design choices we see in industry models. Our contributions are summarized as follows:
  • We open-source more than 4000 checkpoints cumulatively trained on over 10 trillion tokens. The models we provide are diverse across architectural and training hyperparameter axes, enabling more granular studies than previous work (see Figure 2).
  • We highlight the fragility and common pitfalls of prior scaling laws. There are many decisions to make when choosing points to fit scaling laws that significantly change the slope of the law (see Figure 6).
  • We modify standard scaling laws to account for the width and depth of models. Our scaling laws estimate not only parameter count, but also the optimal aspect ratio for any given training budget. We compare our estimates to industrial models (see Figure 5).
  • We modify scaling laws to account for practical considerations like GPU efficiency and overtraining. When distorted to account for training and inference costs, scaling laws predict considerably wider and shallower models (see Figure 7).
  • We consider what happens when scaling laws are inaccurate. Given that scaling laws are quite sensitive to design decisions, it is natural to ask how best to cope with their imperfections. We find that it is much better to err on the side of wider models than narrow ones. We also find that it is better to err on the side of smaller models, as overtraining is quite efficient (see Figure 8).
@@ 2 Related Work
2.1 Scaling Laws
@@
Scaling laws address the trade-off between parameter count and number of training tokens, attempting to find the minimum loss possible for a language model with a constrained FLOP budget. Unfortunately, scaling laws treat model design and training as if it has a single dimension (parameter count). In reality, training is sensitive to many choices. Notably, Hoffmann et al. (2022) find significantly different fitted laws (Equation 1) compared to Kaplan et al. (2020). Pearce & Song (2024) and Porian et al. (2024) attribute most of this discrepancy to the choice to exclude embedding parameters from the parameter count, both showing one law can be transformed into the other via controlled changes. Kaplan et al. (2020) justify including embedding parameters by showing that non-embedding parameters have a cleaner relationship with test loss. Scaling laws are also commonly included in many large model releases (Hu et al., 2024; Bi et al., 2024; Dubey et al., 2024).
@@
Choshen et al. (2024) collect both loss and benchmark performance metrics for a multitude of models and offer a practitioner’s guide to fitting scaling laws. Most notably, they suggest $5$ models are ample to fit a scaling law, and the early period of training should be excluded from the analysis. In contrast, Li et al. (2024b) show that selecting data according to different tokens-per-parameter ratios and using small grid searches when fitting scaling laws can cause big swings in outcomes. Hägele et al. (2024) suggest that a constant learning rate plus cooldown is preferable to a cosine learning rate schedule. The authors also find that stochastic weight averaging should be encouraged in scaling law analysis as it tends to lead to better models. Furthermore, Inbar & Sernau (2024) observe that FLOPs cannot be used to predict wall-clock time nor memory movement, and suggest that fast-training architectures may be preferred over those prescribed by scaling laws.
There are multiple works analyzing whether scaling laws can be used to predict downstream performance. Ruan et al. (2024) show that scaling laws can be predictive of benchmark performance. Caballero et al. (2023) suggest broken scaling laws that predict performance of both downstream and upstream tasks. Works in this vein are myriad; see our extended literature review in Appendix B.
Figure 1: The meaning of width and depth. We visualize a standard transformer architecture, highlighting the “width” as the size of the hidden dimension and the “depth” as the number of transformer blocks. These are the quantities we’re interested in providing a prescription for with our new scaling laws.
2.2 The Role of Model Shape
@@
Levine et al. (2020) find that, for large models, optimal depth grows logarithmically with width. Henighan et al. (2020) find there is an optimal aspect ratio for each modality they study which gives maximum performance, for example they find $5$ to be optimal for math models. Petty et al. (2024) claim small ($<$400M) transformers have diminishing benefits from depth. Brown et al. (2022) show that in some cases shallower models can beat their parameter equivalent deep models on tasks for encoder-decoder transformer architectures. This differs from Kaplan et al. (2020) who suggest aspect ratio is not a determining factor for final loss.
Most similarly to our work, Alabdulmohsin et al. (2024) study the impact of width and depth for encoder-decoder vision transformers, using their laws to create a smaller transformer model which has competitive downstream performance when compared with much larger models. The architecture found in this study has since been used by Beyer et al. (2024) in PaliGemma.
Model shape has also been analyzed for sparse mixture of expert models and in the context of finetuning. Krajewski et al. (2024) use “granularity” to allow their law for mixture of expert models to predict the width of the experts. Tay et al. (2022) show that downstream performance strongly depends on shape when finetuning but pretraining perplexity does not.
2.3 Zero-shot Hyperparameter Transfer
The ability to train a series of models with extremely different parameter counts is an implicit requirement of any scaling law analysis. Work on zero-shot hyperparameter transfer across transformer model widths is mature (Yang et al., 2021; Everett et al., 2024; Hayou & Yang, 2023; Dey et al., 2024), but achieving transfer across diverse model depths is less well studied, especially in transformer language models (Bordelon et al., 2024).
@@ 3 Designing Our Scaling Laws
We detail the design of our scaling laws, including model selection, the choice of learning rate, and curve fitting schemes in the subsequent sections.
@@
To reduce the search space of all possible models we add some constraints, each of which are either based on precedent from a popular model series like Gemma (Team et al., 2024a, b), Llama (Touvron et al., 2023), Pythia (Biderman et al., 2023), or practical considerations such as hardware details.
All models have a head size of $128$ because $256$ is the maximum head dimension supported by the AMD implementation of Flash Attention 2 we utilize and we constrain our search to models with $>1$ attention heads. We assume the simple convention of the Llama series where the head dimension is always the embedding dimension divided by the number of heads, implying that the embedding dimension (width) must be divisible by $128$. Following conventions from the Gemma suite, we constrain the head count to be even to enable Grouped Query Attention (Ainslie et al., 2023) with a query to key ratio of $2:1$ and we fix the intermediate size to be 4×4$\times$4 × the width of the model. We choose our vocabulary size to match the $50,304$ tokens in the Pythia tokenizer. While many of the architecture choices mirror those from Gemma, for simplicity we do not use logit softcapping nor do we tie the embedding and language modeling head weight matrices.
Within these constraints, we search the set of feasible models within target parameter count groups $50M,100M,500M,1B$ and $2B$ with a tolerance of ±$\pm 5\%$. At smaller scales we train up to 5 models at diverse widths and depths. At large parameter counts we train only three models, aiming for one “standard” aspect ratio (similar to existing models), one “wide” model, and one “deep” model. We visualize the models we choose to train in Figure 2 overlaid with a selection of existing models from prior work. In the Appendix we plot the entire discrete set of all possible models under our constraints (Figure 22). Our $22$ different models range from $50M$ to $2B$ parameters, spanning $11$ widths from $256$ to $3072$ and $18$ depths from $3$ to $80$.
@@
For the main set of training runs, we train each model for $350B$ tokens of Dolma (Soldaini et al., 2024) data with a context length of $2048$ and a world batch size of $2048$ sequences. We use a linear learning rate warm up over $80$ million tokens, and then train at a constant learning rate, which we adjust for model size as described in Section 3.1.
@@
We also perform ablations over both cooldown and learning rate. For the cooldown ablation, we take the checkpoints saved every $10$ billion tokens for the the first $100$ billion tokens of training and cool these down creating a second set of models which have had their learning rate annealed to 00 linearly. Specifically, we cool each model down by training for a further $10\%$ of the total tokens which it has seen during training, i.e. our cooled down set of models have training budgets ranging from $11$ to $110$ billion tokens.

### v2

@@ Abstract
Scaling laws are typically fit using a family of models with a narrow range of frozen hyper-parameter choices. In this work we study scaling laws using multiple architectural shapes and hyperparameter choices, highlighting their impact on resulting prescriptions. As a primary artifact of our research, we release the Gemstones: an open-source scaling law dataset, consisting of over 4000 checkpoints from transformers with up to 2 billion parameters and diverse architectural shapes; including ablations over learning rate and cooldown. Our checkpoints enable more complex studies of scaling, such as analyzing the relationship between width and depth. By examining our model suite, we find that the prescriptions of scaling laws can be highly sensitive to the experimental design process and the specific model checkpoints used during fitting.
@@ 1 Introduction
Figure 1: The meaning of width and depth. We visualize a standard transformer architecture, highlighting the “width” as the size of the hidden dimension and the “depth” as the number of transformer blocks.
In this work, we produce a vast array of model checkpoints for studying how model design and model selection impact scaling laws. Our models, called the Gemstones because they are loosely based on scaled-down variants of the Gemma architecture, vary in their parameter count, width/depth ratio, training tokens, learning rates, and cooldown schedules. By fitting scaling laws to these checkpoints, we confirm that scaling law parameters and interpretations indeed depend strongly on the selection of models and fitting procedure used, and we quantify the degree to which these decisions impact predictions. By exploiting the variation among our model checkpoints, we also analyze the impact of architectural shape across loss, benchmark performance and training time with findings consistent with design choices we see in industry models. Our contributions are summarized as follows:
We open-source more than 4000 checkpoints cumulatively trained on over 10 trillion tokens. The models we provide are diverse across architectural and training hyperparameter axes, enabling more granular studies than previous work (see Figure˜2).
We highlight the fragility and common pitfalls of prior scaling laws. There are many decisions to make when choosing points to fit scaling laws that significantly change the slope of the law (see Table˜1).
We analyze the impact of model shape on loss, benchmarks and training time. We find that although deep models achieve lower loss and benchmark error when measured using floating point operations, they require significantly more training time when using typical open-source parallelism frameworks (see Figures 4, 7 and 8).
@@ 2 Related Work
Scaling laws address the trade-off between parameter count and number of training tokens, attempting to find the minimum loss possible for a language model with a constrained floating point operation (FLOP) budget. The body of work on scaling laws is vast. Therefore, while we provide a brief overview of key prior work here to contextualize our contributions, we also include an extended literature review in Appendix˜B.
Design Choices for Scaling Laws
@@
Scaling laws often treat model design and training as if it has a single dimension (parameter count), while, in practice, training is sensitive to many choices. Notably, Hoffmann et al. (2022) find significantly different fitted laws (Equation˜1) compared to Kaplan et al. (2020). Pearce and Song (2024) and Porian et al. (2024) attribute most of this discrepancy to the choice to exclude embedding parameters from the parameter count, both showing one law can be transformed into the other via controlled changes. Kaplan et al. (2020) justify excluding embedding parameters by showing that non-embedding parameters have a cleaner relationship with test loss. Scaling laws are also commonly included in many large model releases (Hu et al., 2024; Bi et al., 2024; Dubey et al., 2024).
@@
Choshen et al. (2024) collect both loss and benchmark performance metrics for a multitude of models and offer a practitioner’s guide to fitting scaling laws. Notably, they suggest that $5$ models are ample to fit a scaling law, and that you should omit the early part of training when fitting, because those early steps don’t follow the same scaling behavior and can skew the results. In contrast, Li et al. (2024c) demonstrate that varying the tokens-per-parameter ratio and relying on limited grid searches when fitting scaling laws can lead to large variations in results. Hägele et al. (2024) suggest that a constant learning rate plus cooldown is preferable to a cosine learning rate schedule as all intermediate checkpoints can be used for fitting. The authors also find that stochastic weight averaging should be encouraged in scaling law analysis as it tends to lead to better models. Furthermore, Inbar and Sernau (2024) observe that FLOPs cannot be used to predict wall-clock time nor memory movement, and suggest that fast-training architectures may be preferred over those prescribed by scaling laws.
There are multiple works analyzing whether scaling laws can be used to predict downstream performance. Bhagia et al. (2024) first predict task-specific loss and then use this to predict performance on the task, using a sigmoidal function to map from loss to accuracy. Gadre et al. (2024) predict top-1 error, fitting a power law function on the perplexity of the model to predict error. Gadre et al. (2024) also look at the impact of overtraining, finding scaling laws that extrapolate with the amount of overtraining. Dey et al. (2023) analyze the trade off of inference FLOPs and training FLOPs using scaling laws to prescribe training configurations that balance training and inference. Unfortunately, both Dey et al. (2023) and Biderman et al. (2023) train on the Pile (Gao et al., 2020) which has since been taken down due to copyright, leaving a gap for a model collection in the open literature.
The Role of Model Shape
@@
Another line of research specifically studies the interplay between model width and model depth; for clarity we visualize our working definitions for these dimensions in Figure˜1. Levine et al. (2020) find that, for large models, optimal depth grows logarithmically with width. Henighan et al. (2020) find there is an optimal aspect ratio for each modality they study which gives maximum performance: for example, they find $5$ to be optimal for math models. Team et al. (2024b) compare two 9 billion parameter models and find the deeper model outperforms the wider one consistently across benchmarks. Unfortunately, the authors are vague about the specific details of this result. Petty et al. (2024) claim small (<<400M) transformers have diminishing benefits from depth. Brown et al. (2022) show that in some cases shallower models can beat their parameter-equivalent deep models on tasks for encoder-decoder transformer architectures. These results differ from Kaplan et al. (2020) who suggest aspect ratio is not a determining factor for final loss. Tay et al. (2022) show that downstream performance strongly depends on shape when finetuning but pretraining perplexity does not. Alabdulmohsin et al. (2024) study the impact of width and depth for encoder-decoder vision transformers, using their laws to create a smaller transformer model which has competitive downstream performance when compared with much larger models. The architecture found in this study has since been used by Beyer et al. (2024) in PaliGemma. Concurrently, Zuo et al. (2025) study the impact of width and depth in hybrid architectures, finding that a deeper 1.5B model can match or even outperform 3B and 7B models.
As discussed above, the literature on how the aspect ratio of a LLM affects its performance and scaling characteristics is simultaneously extensive but somewhat inconclusive. While we do not presume to fully answer every question in this space, the experiments we describe in the rest of this work make progress on how to understand the results of prior studies and the impacts of certain architecture choices in a fully open, reproducible, and extensible way.
@@ 3 Designing Our Scaling Laws
We discuss the design of our scaling laws, including model selection, the choice of learning rate, and curve fitting schemes in the subsequent sections and in greater detail in Appendix˜A.
@@
To reduce the search space of all possible models, we add some constraints, each of which are either based on precedent from a popular model series like Gemma (Team et al., 2024a, b), Llama (Touvron et al., 2023), Pythia (Biderman et al., 2023), or practical considerations such as hardware details (see Appendix˜A).
Figure 2: Distribution of prior scaling law models, industry models, and our models in terms of width and depth. Prior work (purple and green) and industry models (blue and orange) mostly lie on a fixed width-depth line.
Within these constraints, we search the set of feasible models within target parameter count groups $50M,100M,500M,1B$ and $2B$ with a tolerance of $\pm 5\%$. At smaller scales we train up to 5 models at diverse widths and depths. At large parameter counts we train only 3 models, aiming for one “standard” aspect ratio (similar to existing models), one “wide” model, and one “deep” model. We visualize the models we choose to train in Figure˜2 overlaid with a selection of existing models from prior work. In the Appendix, we plot the entire discrete set of all possible models under our constraints (Figure˜10). Our $22$ different models range from $50M$ to $2B$ parameters, spanning $11$ widths from $256$ to $3072$ and $18$ depths from $3$ to $80$.
@@
For the main set of training runs, we train each model for $350B$ tokens of Dolma 1.7 (Soldaini et al., 2024) data. We target a total batch size of 4~4 million tokens following(Touvron et al., 2023; Dubey et al., 2024; Bai et al., 2023), with a context length of $2048$ and a world batch size of $2048$ sequences. Following Hägele et al. (2024) and Hu et al. (2024), we use a linear learning rate warm up over $80$ million tokens, and then train at a constant learning rate, which we adjust for model size as described in Section˜A.1.
@@
We perform ablations over both cooldown and learning rate. For the cooldown ablation, we take the checkpoints saved every $10$ billion tokens for the the first $100$ billion tokens of training and cool these down, creating a second set of models which have had their learning rate annealed to 0 linearly. Specifically, we cool each model down by training for a further $10\%$ of the total tokens which it has seen during training, i.e. our cooled down set of models have training budgets ranging from $11$ to $110$ billion tokens. We also ablate our choice of learning rate by running all models for $100$ billion tokens with half of the learning rate we use for our main analysis. The full details of the scalable learning rate and parameter initialization scheme–designed to enable hyperparameter transfer across model sizes and aspect ratios–are provided in Section˜A.1.